S = 5NT - TS + NS
Solve for T

Answer by scott8148(6628)

subtracting NS ___ S - NS = 5NT - TS

factoring ___ S - NS = T (5N - S)

dividing by (5N - S) ___ (S - NS) / (5N - S) = T